OpenStudy (anonymous):

dimensional analysis 400in^3 to ft^3 12m^3 to cm^3 0.4ft^3 to mm^3

OpenStudy (anonymous):

^ will help alot already does all the equations for you.

OpenStudy (anonymous):

@BVB4rmy @robtobey I still need the process. Thanks for helping me!

OpenStudy (anonymous):

in^3 x = ft^3 x= ft^3/in^3=1/12^3

OpenStudy (anonymous):

@robtobey stills a little bit confusing

OpenStudy (anonymous):

found the first one 400in^3= 400in^3*.00057ft^3/1in^3= 0.228ft^3

OpenStudy (anonymous):

Refer to the Mathematica 9 attachment.